# Exploring Multidimensional Checkworthiness: Designing AI-assisted Claim Prioritization for Human Fact-checkers

### v1

@@ Abstract.
Given the massive volume of potentially false claims circulating online, claim prioritization is essential in allocating limited human resources available for fact-checking. In this study, we perceive claim prioritization as an information retrieval (IR) task: just as multidimensional IR relevance, with many factors influencing which search results a user deems relevant, checkworthiness is also multi-faceted, subjective, and even personal, with many factors influencing how fact-checkers triage and select which claims to check. Our study investigates both the multidimensional nature of checkworthiness and effective tool support to assist fact-checkers in claim prioritization. Methodologically, we pursue Research through Design combined with mixed-method evaluation. We develop an AI-assisted claim prioritization prototype as a probe to explore how fact-checkers use multidimensional checkworthiness factors in claim prioritization, simultaneously probing fact-checker needs while also exploring the design space to meet those needs.
Our study with 16 professional fact-checkers investigates: 1) how participants assessed the relative importance of different checkworthy dimensions and apply different priorities in claim selection; 2) how they created customized GPT-based search filters and the corresponding benefits and limitations; and 3) their overall user experiences with our prototype. Our work makes a conceptual contribution between multidimensional IR relevance and fact-checking checkworthiness, with findings demonstrating the value of corresponding tooling support. Specifically, we uncovered a hierarchical prioritization strategy fact-checkers implicitly use, revealing an underexplored aspect of their workflow, with actionable design recommendations for improving claim triage across multi-dimensional checkworthiness and tailoring this process with LLM integration. Overall, our findings provide insights into fact-checker work practices and design implications for more personalized and efficient claim prioritization.
@@ 1. Introduction
Given the need for human fact-checking, claim prioritization is key to efficiently allocating human resources (Westlund et al., 2022; Micallef et al., 2022; Sehat et al., 2024). Prioritization seeks to triage and select the most consequential claims by considering checkworthiness factors aligned with the goals and news values of fact-checking organizations (Vinhas and Bastos, 2023; Bélair-Gagnon et al., 2023). While many NLP methods have been developed to identify and monitor misinformation, NLP research has typically sought to automate rather than develop mixed-initiative (Horvitz, 1999) tools to assist human fact-checkers in claim prioritization. Given the complexity and uncertainty underlying assessment of checkworthiness (Micallef et al., 2022; Procter et al., 2023; Liu et al., 2024), better tooling could significantly help.
@@
Our prototype provides two key capabilities, beyond a basic search box for entering query terms. First, we provide automated models that predict four checkworthiness dimensions (“Verifiable”, “Likely harmful”, “Likely false”, and “Interest to the public”), coupled with simple UI slider widgets that support dynamically varying the relative weight assigned to each dimension of checkworthiness, individually or in combination, to customize claim ranking in real-time. The second key capability enables fact-checkers to develop additional zero-code, custom search filters using large language model (LLM) technology. This allows additional dimensions of checkworthiness to be introduced and influence claim ranking beyond the four dimensions supported natively. More generally, the flexibility and power of the customized LLM search filter can help fact-checkers to overcome limitations of traditional keyword search. Just as LLMs enable non-programmers to quickly formulate new AI tasks without model training and data acquisition, our goal was to enable fact-checkers to specify custom search criteria for claim prioritization in natural language.
Our user study with 16 professional fact-checkers employed mixed-method evaluations to collect and analyze participant experiences and reflections. Guided by an RtD process, our prototype enabled us to probe and observe how fact-checkers flexibly triage claims across multidimensional checkworthiness. We investigate: 1) how participants assessed the relative importance of different checkworthy dimensions and applied different priorities in claim selection; 2) how they created customized LLM-based search filters and the corresponding benefits and limitations; and 3) their overall user experiences with our prototype.
Research Contributions: In this paper, we examine how fact-checkers dynamically prioritize claims by considering various checkworthiness factors and uncover specific user needs for tools to support this process. Specifically, we developed an interactive claim prioritization prototype as a design probe to investigate these dynamics in-depth. We uncovered a hierarchical prioritization strategy that they implicitly use, shedding light on an underexplored aspect of their workflow. We also synthesized actionable design recommendations learned from fact-checkers, suggesting mechanisms to better triage claims across multi-dimensional checkworthiness and to tailor this process by integrating LLMs. These insights deepen our understanding of fact-checker work practices and the supporting tools they require, while also offering broader design implications for improving relevance judgment and triage in other user activities.
@@ 2.1. Journalistic Fact-checking and Digital Tools Used for Claim Prioritization
Fact-checkers use many tools to search for, monitor, filter, and collect potential claims to check. As reported in prior work (Micallef et al., 2022; Beers et al., 2020; Westlund et al., 2022; Procter et al., 2023; Das et al., 2023), off-the-shelf tools include Google search, search provided by the social media platforms, third-party monitoring software, for example, TweetDeck [^1] TweetDeck has now become X Pro: https://pro.twitter.com/ and CrowdTangle [^2] https://www.crowdtangle.com/, and other open-source intelligent tools. More advanced tools were also built to better identify checkable claims that contain factual statements. For example, Majithia et al. (2019) build ClaimPortal, a tool incorporating Claimbuster (Hassan et al., 2017), to identify checkable claims from Tweets and perform traffic analytics. The UK fact-checking organization, Full Fact [^3] https://fullfact.org/ai/about/, built a claim monitoring tool that helps identify different types of claims from different media and news outlets, including statistical, opinionated, and predicted claims. Meta also provides fact-checking tools that enable fact-checking organizations they partnered with [^4] https://transparency.meta.com/features/how-fact-checking-works to monitor, search, and check claims on their platforms, including Facebook and Instagram.
@@ 2.2. Relevance Judgment and Claim Checkworthiness Assessment
Since fact-checkers primarily prioritize claims within the context of information search, we frame their process of searching for and selecting claims as analogous to a traditional Information Retrieval (IR) task, where a user has an information need (Wilson, 1981) and seeks relevant information to address that need. Just as relevance is multidimensional, with many factors influencing which search results are deemed relevant (Xu and Chen, 2006), checkworthiness is also multi-faceted affected by diverse factors (Alam et al., 2021). Given this parallel, we review key literature from both IR and fact-checking to ground our understanding of claim prioritization and to inform our tool design.
In IR studies, evaluating relevance through multiple dimensions provides a more accurate assessment of search results than considering only unidimensional topical relevance. Jiang et al. (2017) integrates four factors—“Novelty,” “Understandability,” “Reliability,” and “Effort”—with user experience measures to evaluate an IR system. This multidimensional approach to judging relevance showed a stronger statistical relationship with user experience than just topical relevance. It is also recognized that individuals perceive the importance of these multidimensional factors differently, which can influence their final judgment of search relevance. Zhang et al. (2014) use structural equation modeling to examine the relationship between five-dimensional factors and the overall relevance of search results. The findings from their model indicated a considerable difference in the weight individuals assign to different relevance factors. Additional research shows that user perceptions of multidimensional relevance are further influenced by user domain expertise (Tamine and Chouquet, 2017) and biases (Krieg et al., 2022). This suggests that relevance judgment is highly subjective and individualized.
@@
NLP research has also annotated claims for different checkworthy dimensions in order to build and test predictive models (Hassan et al., 2017; Konstantinovskiy et al., 2021; Alam et al., 2021), with annotation guidelines using linguistic features to infer various aspects of checkworthiness (Allein and Moens, 2020). A well-known NLP competition, CLEF CheckThat! (Barrón-Cedeño et al., 2024, 2023), emphasizes the multidimensional nature of evaluating claim checkworthiness and helps capture some of them from human annotators. Despite the prominence of this work, there is little insight into provenance of annotated dimensions: how the particular dimensions chosen were selected or their relative importance, what other dimensions might exist, how fact-checkers use these dimensions in practice, etc.
Informed by decades of study of relevance in IR (Saracevic, 1975, 2007a, 2007b), we can infer that fact-checkers perceive the importance of various dimensions of checkworthiness differently, which influences how they identify and select claims during searches. Therefore, several open questions arise that exemplify the gaps in existing claim prioritization: How do fact-checkers evaluate the relative importance of various dimensions? Are different dimensions of checkworthiness considered serially or in parallel by fact-checkers to conduct claim selection effectively? How do fact-checker self-reported perspectives on these dimensions align with their actual behavior during claim search and selection? When different dimensions compete in claim ranking, how do fact-checkers navigate these trade-offs dynamically? In Section 3.4.4, we structure these questions around the primary research goal and explain how we address them. As we seek to address some of these questions with an RtD approach, we synthesize important literature on RtD in the next section and describe our motivation of using RtD.
@@ 2.3. Research through Design (RtD) in Misinformation Research
As discussed in the end of Section 2.2, our study investigates how fact-checkers perceive the relative importance of checkworthiness dimensions and apply this in claim prioritization during search, to better reveal potential user needs. This knowledge is situational and contextualized within fact-checking, especially in fact-checker behavior of searching, filtering, and selecting claims. RtD is a well-suited approach for addressing this challenge. It emphasizes using design as a method to uncover nuanced, situational knowledge rather than focusing on developing definitive solutions, particularly when user practices are not yet fully understood (Gaver, 2012; Zimmerman et al., 2010). In our context, this involves exploring how fact-checkers prioritize various dimensions of checkworthiness, how they apply this situational understanding to the process of searching for claims, and what AI solutions we can design to facilitate this process more effectively.
@@ 3. Research through Design
In Zimmerman et al. (2010)’s examinations of RtD practices across different design projects, they found that “All of the projects employed a RtD approach, creating artifacts that included products, prototypes and models that illustrated future visions, uses of new materials, and potential ideas.” Bowers (2012) also emphasize that the artifacts created via RtD aims to “provide the design research community with information about how to design.” Thus, RtD differs form traditional empirical studies, such as interviews or focus groups, where these methods typically do not involve user interactions with artifacts. For understanding user practices, these methods primarily rely on verbal accounts and thematic analysis of participant narratives, rather than gaining insights through direct observations of participant actions and processes. It also contrasts with tool evaluations, where the study focus is on assessing a tool usability rather than better uncovering user practices and generating design knowledge to inform future solutions (Frankel and Racine, 2010; Pierce, 2014).
@@ 3.1. Frame Research Goals
After interviewing fact-checkers around the globe, Sehat et al. (2024) reported “no established systematic approach towards claim prioritization” today. Additionally, fact-checker decisions to focus on specific claims typically depend on case-by-case situations and are heavily influenced by the local news context (Liu et al., 2024; Procter et al., 2023; Micallef et al., 2022). Thus, both theoretical and empirical understanding highlights the complexity and uncertainty involved when fact-checkers use tools to search and select claims to check. This poses a further design consideration: new tools developed for claim prioritization should support such complexity and uncertainty with necessary flexibility to adapt to dynamically changing priorities. While prior work has explored aspects of this challenge, it often stops short of offering clear solutions in tool supports. This gap presents an opportunity to leverage RtD to better inform and develop actionable design outcomes.
In particular, we believe using artifact-driven actions and reflections is more effective to first untangle the complexity and uncertainty of fact-checker tasks in claim prioritization. As noted by Bowers (2012) and Pierce (2014), the value of using RtD lies in creating exploratory artifacts to uncover design knowledge. In our context, these artifacts act as probes, enabling fact-checkers to demonstrate how they dynamically triage claims across various checkworthiness factors. By observing the strategies they develop and adapt over time to improve the efficiency of searching and selecting claims, we could gain valuable insights into user behaviors and patterns, which subsequently guide the design of future tools.
@@
    A practice-based examination of fact-checker practice and needs for claim prioritization. We aim to build a prototype as a probe to elicit fact-checker insights into how they flexibly triage claims among multidimensional checkworthiness by searching, filtering, and selecting claims in real time. Additionally, as RtD contributes to the creation of innovative solutions (Gaver, 2012; Zimmerman and Forlizzi, 2014), we also synthesize insights, including user behaviors and patterns into more sophisticated needs and concrete design suggestions for claim prioritization based on what we observed when fact-checkers use the prototype.
@@
3.2. Discover and Define Design Choices
@@
To synthesize design work from both academic research and industry practices for our tool-building, our literature search focused on three main aspects: 1) existing tools used by fact-checkers to identify and monitor claims; 2) traditional multi-faceted search, browsing, and filtering that support general information seeking; and 3) customizable functions that provide users with a more personalized search experience. We conducted an academic literature search on Google Scholar using keywords such as “multi-faceted search,” “personalized search,” and “browsing and filtering.” Most of the academic research we identified originated from conferences related to SIGCHI, SIGIR, and CHIIR. To explore tools currently used by fact-checkers, we examined a range of existing resources and tools related to claim prioritization. These included a collection of tools curated by the nonprofit research organizations, such as RAND Corporation and Credibility Coalition.
@@
We summarize our search results in Table 2. These include standard keyword and semantic search, multi-faceted filters, personalized weighting, and user-generated facets. These features illustrate the different levels of customization that assist users in their search and browsing activities. We make a simplifying assumption in restricting our scope to textual claims. While future work should investigate the multi-modal setting as well, we show that even this simplified setting is sufficient to reveal broadly useful insights into fact-checker needs and corresponding design implications.
Categories Features Description Academic or industry implementation
@@
Search Keyword search Users enter keywords to look for exact matches of documents where the keyword appears. (Yee et al., 2003; Jasim et al., 2022; Majithia et al., 2019; Guo et al., 2023) Google fact-check tools, Trendolizer, Meta fact-checking tool, Full Fact Alpha
@@
Semantic search Users enter a search query to retrieve documents that provide contextual meanings similar to the query. (Jasim et al., 2022; Majithia et al., 2019) Google fact-check tools, Trendolizer, Meta fact-checking tool, Full Fact Alpha
@@
Image reverse search Users use an image as a search query to find similar images from the database. (Cai et al., 2019) Google fact-check tools
@@
Filtering Multi-faceted filters Users select various criteria from different categories to dynamically filter documents. Only documents relevant to the criteria are updated. (Yee et al., 2003; Kern et al., 2018; Cai et al., 2019; Jasim et al., 2022; Majithia et al., 2019; Guo et al., 2023) Trendolizer, Meta fact-checking tool, Full Fact Alpha
@@
Personalized weighting Users adjust the importance of certain criteria or elements to tailor the retrieved results based on individual preferences. (Kern et al., 2018; Cai et al., 2019)
@@
User-generated facets Users create and define the criteria or categories used for organizing and filtering research results. (Kern et al., 2018; Papenmeier et al., 2023)
@@
For example, at this design stage, we first decided to develop a multi-dimensional ranking system in order to enhance algorithmic transparency, enabling fact-checkers to explore how different facets of checkworthiness could influence a unified, overall checkworthiness ranking. We also know from prior work in other expert search domains (e.g., legal (Paul and Baron, 2006; Oard, 2013) and medical search (Wallace et al., 2013)) that experts highly value transparent, controllable ranking functions. Our tool seeks to provide such transparent ranking by: 1) selecting established dimensions of checkworthiness; 2) making those dimensions explicit, visible, and actionable in the user interface, and 3) providing real-time updates to search results in response to user weight changes. Our work thus builds on both prior technical work as well as prior understanding of user needs in expert search domains, extending these capabilities to support fact-checking tasks today.
In addition, because fact-checkers (individually or organizationally) may value different aspects of checkworthiness that were not captured in our standard facets, we further integrated LLM-customization as another feature, allowing fact-checkers to extend our tool by incorporating additional checkworthiness. Relatedly, we also expected that algorithmic transparency would be promoted by keeping the search bar (topical relevance) separate from checkworthiness, motivating separation between search bar vs. the standard and LLM-customized checkworthiness facets.
More generally, we recognized that separating these functionalities aligns more effectively with our research goals. As RtD prioritizes the discovery of design knowledge over the pursuit of optimal solutions, particularly in contexts where user practices are ambiguous (Zimmerman and Forlizzi, 2014; Gaver, 2012; Frankel and Racine, 2010) (in our case i.e., claim prioritization), we viewed these three features – topical search bar, standard checkworthiness filters, and LLM-customized filters – as distinct yet valuable tools. They enabled participants to articulate the differences and uncover use-inspired insights, guiding decisions about whether to merge these features together or maintain them as separate components for designing future tools.

### v2

@@ Abstract.
Given the massive volume of potentially false claims circulating online, claim prioritization is essential in allocating limited human resources available for fact-checking. In this study, we perceive claim prioritization as an information retrieval (IR) task: just as multidimensional IR relevance, with many factors influencing which search results a user deems relevant, checkworthiness is also multi-faceted, subjective, and even personal, with many factors influencing how fact-checkers triage and select which claims to check. Our study investigated both the multidimensional nature of checkworthiness and effective tool support to assist fact-checkers in claim prioritization. Methodologically, we pursued Research through Design combined with mixed-method evaluation.
Specifically, we developed an AI-assisted claim prioritization prototype as a probe to explore how fact-checkers use multidimensional checkworthy factors to prioritize claims, simultaneously probing fact-checker needs and exploring the design space to meet those needs. With 16 professional fact-checkers participating in our study, we uncovered a hierarchical prioritization strategy fact-checkers implicitly use, revealing an underexplored aspect of their workflow, with actionable design recommendations for improving claim triage across multidimensional checkworthiness and tailoring this process with LLM integration.
@@ 1. Introduction
Given the need for human fact-checking, claim prioritization is key to efficiently allocating human resources (Westlund et al., 2022; Micallef et al., 2022; Sehat et al., 2024). Prioritization seeks to triage and select the most consequential claims by considering checkworthiness factors aligned with the goals and news values of fact-checking organizations (Vinhas and Bastos, 2023; Bélair-Gagnon et al., 2023). While many NLP methods have been developed to identify and monitor misinformation, NLP research has typically sought to automate rather than develop mixed-initiative (Horvitz, 1999) tools to assist human fact-checkers in claim prioritization. Given the complexity and uncertainty underlying the assessment of checkworthiness (Micallef et al., 2022; Procter et al., 2023; Liu et al., 2024), better tooling could significantly help.
@@
Our prototype provides two key capabilities beyond a basic search box for entering query terms. First, we provide automated models that predict four checkworthiness dimensions (“Verifiable”, “Likely harmful”, “Likely false”, and “Interest to the public”), coupled with simple UI slider widgets that support dynamically varying the relative weight assigned to each dimension of checkworthiness, individually or in combination, to customize claim ranking in real-time. The second key capability enables fact-checkers to develop additional zero-code, custom search filters using large language model (LLM) technology. This allows additional dimensions of checkworthiness to be introduced and influence claim ranking beyond the four dimensions supported natively. More generally, the flexibility and power of the customized LLM search filter can help fact-checkers to overcome the limitations of traditional keyword search. Just as LLMs enable non-programmers to quickly formulate new AI tasks without model training and data acquisition, our goal was to enable fact-checkers to specify custom search criteria for claim prioritization in natural language.
Our user study with 16 professional fact-checkers employed mixed-method evaluations to collect and analyze participant experiences and reflections. Guided by an RtD process, the prototype enabled us to probe and observe how fact-checkers flexibly triage claims across multidimensional checkworthiness. We investigate: 1) how participants assessed the relative importance of different checkworthy dimensions and developed priorities in claim selection; 2) how they created customized LLM-based search filters and the corresponding benefits and limitations; and 3) their overall user experiences with our prototype.
Research Contributions: In this paper, we examine how fact-checkers dynamically prioritize claims by considering various checkworthiness factors and uncover specific user needs for tools to support this process. Specifically, we developed an interactive claim prioritization prototype as a design probe to investigate these dynamics in-depth. We uncovered a hierarchical prioritization strategy that they implicitly use, shedding light on an underexplored aspect of their workflow. We also synthesized actionable design recommendations learned from fact-checkers, suggesting mechanisms to better triage claims across multidimensional checkworthiness and to tailor this process by integrating LLMs. These insights deepen our understanding of fact-checker work practices and the supporting tools they require, while also offering broader design implications for improving relevance judgment and triage in other user activities.
@@ 2.1. Journalistic Fact-checking and Digital Tools Used for Claim Prioritization
Fact-checkers use many tools to search for, monitor, filter, and collect potential claims to check. As reported in prior work (Micallef et al., 2022; Beers et al., 2020; Westlund et al., 2022; Procter et al., 2023; Das et al., 2023), off-the-shelf tools include Google search, search provided by the social media platforms, third-party monitoring software [^1] TweetDeck has now become X Pro: https://pro.twitter.com/, CrowdTangle (suspended): https://www.crowdtangle.com/., for example, TweetDeck and CrowdTangle, and other open-source intelligent tools. More advanced tools were also built to better identify checkable claims that contain factual statements. For example, Majithia et al. (2019) build ClaimPortal, a tool incorporating Claimbuster (Hassan et al., 2017), to identify checkable claims from Tweets and perform traffic analytics. The UK fact-checking organization, Full Fact [^2] Full Fact AI: https://fullfact.org/ai/about/., built a claim monitoring tool that helps identify different types of claims from different media and news outlets, including statistical, opinionated, and predicted claims. Meta also provided fact-checking tools [^3] Meta’s fact-checking partnership (suspended at the time of publication): https://transparency.meta.com/features/how-fact-checking-works. that enabled fact-checking organizations they partnered with to monitor, search, and check claims on their platforms, including Facebook and Instagram.
@@ 2.2. Relevance Judgment and Claim Checkworthiness Assessment
Since fact-checkers primarily prioritize claims within the context of information search, we perceive their process of searching for and selecting claims as analogous to a traditional Information Retrieval (IR) task, where a user has an information need (Wilson, 1981) and seeks relevant information to address that need. Just as relevance is multidimensional, with many factors influencing which search results are deemed relevant (Xu and Chen, 2006), checkworthiness is also multi-faceted and affected by diverse factors (Alam et al., 2021). Given this parallel, we review key literature from both IR and fact-checking to ground our understanding of claim prioritization and to inform our tool design.
In IR studies, evaluating relevance through multiple dimensions provides a more accurate assessment of search results than considering only unidimensional topical relevance. Jiang et al. (2017) integrates four factors—“Novelty,” “Understandability,” “Reliability,” and “Effort”—with user experience measures to evaluate an IR system. This multidimensional approach to judging relevance showed a stronger statistical relationship with user experience than just topical relevance. It is also recognized that individuals perceive the importance of these multidimensional factors differently, which can influence their final judgment of search relevance. Zhang et al. (2014) use structural equation modeling to examine the relationship between five-dimensional factors and the overall relevance of search results. Findings from their model indicated a considerable difference in the weight that individuals assigned to different relevance factors. Additional research shows that user perceptions of multidimensional relevance are further influenced by user domain expertise (Tamine and Chouquet, 2017) and biases (Krieg et al., 2022). This suggests that relevance judgment is highly subjective and individualized.
@@
NLP research has also annotated claims for different checkworthy dimensions in order to build and test predictive models (Hassan et al., 2017; Konstantinovskiy et al., 2021; Alam et al., 2021), with annotation guidelines using linguistic features to infer various aspects of checkworthiness (Allein and Moens, 2020). A well-known NLP competition, CLEF CheckThat! (Barrón-Cedeño et al., 2024, 2023), emphasizes the multidimensional nature of evaluating claim checkworthiness and helps capture some of them from human annotators. Despite the prominence of this work, there is little insight into the provenance of annotated dimensions: how the particular dimensions were selected or their relative importance, what other dimensions might exist, how fact-checkers used these dimensions in practice, etc.
Informed by decades of study of relevance in IR (Saracevic, 1975, 2007a, 2007b), we can infer that fact-checkers perceive the importance of various dimensions of checkworthiness differently, which influences how they identify and select claims during searches. Therefore, several open questions arise that exemplify the gaps in existing claim prioritization: How do fact-checkers evaluate the relative importance of various dimensions? Are different dimensions of checkworthiness considered serially or in parallel by fact-checkers to conduct claim selection effectively? How do fact-checker self-reported perspectives on these dimensions align with their actual behavior during claim search and selection? When different dimensions compete in claim ranking, how do fact-checkers navigate these trade-offs dynamically? In Section 3.4.4, we structure these questions around the primary research goal and explain how we address them. As we seek to address some of these questions with an RtD approach, we synthesize important literature on RtD in the next section and describe our motivation for using RtD.
@@ 2.3. Research through Design (RtD) in Misinformation Research
As discussed at the end of Section 2.2, our study investigates how fact-checkers perceive the relative importance of checkworthiness dimensions and apply this in claim prioritization during search to better reveal potential user needs. This knowledge is situational and contextualized within fact-checking, especially in fact-checker behavior of searching, filtering, and selecting claims. RtD is a well-suited approach for addressing this challenge. It emphasizes using design as a method to uncover nuanced, situational knowledge rather than focusing on developing definitive solutions, particularly when user practices are not yet fully understood (Gaver, 2012; Zimmerman et al., 2010). In our context, this involves exploring how fact-checkers prioritize various dimensions of checkworthiness, how they apply this situational understanding to the process of searching for claims, and what AI solutions we can design to facilitate this process more effectively.
@@ 3. Research through Design
In Zimmerman et al. (2010)’s examinations of RtD practices across different design projects, they found that “All of the projects employed an RtD approach, creating artifacts that included products, prototypes, and models that illustrated future visions, uses of new materials, and potential ideas.” Bowers (2012) also emphasize that the artifacts created via RtD aim to “provide the design research community with information about how to design.” Thus, RtD differs from traditional empirical studies, such as interviews or focus groups, where these methods typically do not involve user interactions with artifacts. To understand user practices, these methods primarily rely on verbal accounts and thematic analysis of participant narratives, rather than gaining insights through direct observations of participant actions and processes. RtD also contrasts with tool evaluations, where the study focuses on assessing a tool usability rather than better uncovering user practices and generating design knowledge to inform future solutions (Frankel and Racine, 2010; Pierce, 2014).
@@ 3.1. Frame Research Goals
After interviewing fact-checkers around the globe, Sehat et al. (2024) reported “no established systematic approach towards claim prioritization” today. Additionally, fact-checker decisions to focus on specific claims typically depend on case-by-case situations and are heavily influenced by the local news context (Liu et al., 2024; Procter et al., 2023; Micallef et al., 2022). Thus, both theoretical and empirical understanding highlight the complexity and uncertainty involved when fact-checkers use tools to search and select claims to check. This poses a further design consideration: new tools developed for claim prioritization should support such complexity and uncertainty with the necessary flexibility to adapt to dynamically changing priorities. While prior work has explored aspects of this challenge, it often stops short of offering clear solutions in tool support. This gap presents an opportunity to leverage RtD to better inform and develop actionable design outcomes.
In particular, we believe using artifact-driven actions and reflections is more effective in first untangling the complexity and uncertainty of fact-checker tasks in claim prioritization. As noted by Bowers (2012) and Pierce (2014), the value of using RtD lies in creating exploratory artifacts to uncover design knowledge. In our context, these artifacts act as probes, enabling fact-checkers to demonstrate how they dynamically triage claims across various checkworthiness factors. By observing the strategies they develop and adapt over time to improve the efficiency of searching and selecting claims, we could gain valuable insights into user behaviors and patterns, which subsequently guide the design of future tools.
@@
    A practice-based examination of fact-checker practice and needs for claim prioritization. We aim to build a prototype as a probe to elicit fact-checker insights into how they flexibly triage claims among multidimensional checkworthiness by searching, filtering, and selecting claims in real time. Additionally, as RtD contributes to the creation of innovative solutions (Gaver, 2012; Zimmerman and Forlizzi, 2014), we also synthesize insights, including user behaviors and patterns, into more sophisticated needs, as well as concrete design suggestions for claim prioritization based on what we observed when fact-checkers use the prototype.
@@
3.2. Discover and Define Design Probe
@@
To synthesize design work from both academic research and industry practices for our tool-building, our literature search focused on three main aspects: 1) existing tools used by fact-checkers to identify and monitor claims; 2) traditional multi-faceted search, browsing, and filtering that support general information seeking; and 3) customizable functions that provide users with a more personalized search experience. We conducted an academic literature search on Google Scholar using keywords such as “multi-faceted search,” “personalized search,” and “browsing and filtering.” Most of the academic research we identified originated from conferences related to SIGCHI, SIGIR, and CHIIR. To explore tools currently used by fact-checkers, we examined a range of existing resources and tools related to claim prioritization. These included a collection of tools curated by nonprofit research organizations, such as RAND Corporation [^4] RAND Corporation: https://www.rand.org/research/projects/truth-decay/fighting-disinformation/search.html and Credibility Coalition [^5] Credibility Coalition: https://credibilitycoalition.org/.
@@
We summarize our search results in Table 2. These include standard keyword and semantic search, multi-faceted filters, personalized weighting, and user-generated facets. These features illustrate the different levels of customization that assist users in their search and browsing activities. We make a simplifying assumption by restricting our scope to textual claims. While future work should investigate the multi-modal setting as well, we show that even this simplified setting is sufficient to reveal broadly useful insights into fact-checker needs and corresponding design implications.
Categories Features Description Tool implementations
@@
Search Keyword search Users enter keywords to look for exact matches of documents where the keyword appears. Academic: (Yee et al., 2003; Jasim et al., 2022; Majithia et al., 2019; Guo et al., 2023) Industrial: Google fact-check tools, Trendolizer, Meta fact-checking tool, Full Fact Alpha
@@
Semantic search Users enter a search query to retrieve documents that provide contextual meanings similar to the query. Academic: (Jasim et al., 2022; Majithia et al., 2019) Industrial: Google fact-check tools, Trendolizer, Meta fact-checking tool, Full Fact Alpha
@@
Image reverse search Users use an image as a search query to find similar images from the database. Academic: (Cai et al., 2019) Industrial: Google fact-check tools
@@
Filtering Multi-faceted filters Users select various criteria from different categories to dynamically filter documents. Only documents relevant to the criteria are updated. Academic: (Yee et al., 2003; Kern et al., 2018; Cai et al., 2019; Jasim et al., 2022; Majithia et al., 2019; Guo et al., 2023) Industrial: Trendolizer, Meta fact-checking tool, Full Fact Alpha
@@
Personalized weighting Users adjust the importance of certain criteria or elements to tailor the retrieved results based on individual preferences. Academic: (Kern et al., 2018; Cai et al., 2019)
@@
User-generated facets Users create and define the criteria or categories used for organizing and filtering research results. Academic: (Kern et al., 2018; Papenmeier et al., 2023)
@@
For example, at this design stage, we first decided to develop a multi-dimensional ranking system in order to enhance algorithmic transparency, enabling fact-checkers to explore how different facets of checkworthiness could influence a unified, overall checkworthiness ranking. We also know from prior work in other expert search domains (e.g., legal (Paul and Baron, 2006; Oard, 2013) and medical search (Wallace et al., 2013)) that experts highly value transparent, controllable ranking functions. Our tool seeks to provide such transparent ranking by 1) selecting established dimensions of checkworthiness, 2) making those dimensions explicit, visible, and actionable in the user interface, and 3) providing real-time updates to search results in response to user weight changes. Our work thus builds on both prior technical work and a prior understanding of user needs in expert search domains, extending these capabilities to support fact-checking tasks today.
In addition, because fact-checkers (individually or organizationally) may value different aspects of checkworthiness that were not captured in our standard facets, we further integrated LLM customization as another feature, allowing fact-checkers to extend our tool by incorporating additional checkworthiness. Relatedly, we also expected that algorithmic transparency would be promoted by keeping the search bar (topical relevance) separate from checkworthiness, motivating separation between the search bar vs. the standard and LLM-customized checkworthiness facets.
More generally, we recognized that separating these functionalities aligns more effectively with our research goals. As RtD prioritizes the discovery of design knowledge over the pursuit of optimal solutions, particularly in contexts where user practices are ambiguous (Zimmerman and Forlizzi, 2014; Gaver, 2012; Frankel and Racine, 2010) (in our case, i.e., claim prioritization), we viewed these three features – topical search bar, standard checkworthiness filters, and LLM-customized filters – as distinct yet valuable tools. They enabled participants to articulate the differences and uncover use-inspired insights, guiding decisions about whether to merge these features together or maintain them as separate components for designing future tools.